OpenStudy (anonymous):

The rearrangement of methyl isonitrile (CH3NC) to acetonitrile (CH3CN) is a first-order reaction and has a rate constant of 5.11×10−5s−1 at 472 K. If the initial concentration of CH3NC is 3.30×10−2 M . Part A: What is the molarity of CH3NC after 1.80 h ? Part B: How many minutes does it take for the CH3NC concentration to drop to 2.90×10−2 M ? Part C: How many minutes does it take for 25% of the CH3NC to react?

OpenStudy (cuanchi):

Part A: ln[A]−ln[A]o=−kt => ln[A]=−kt +ln[A]o [A]= e^(−kt +ln[A]o) Part B: ln[A]−ln[A]o=−kt => t = -(ln[A]−ln[A]o)/k Part C: t = -(ln[A]−ln[A]o)/k t = -(ln 75−ln 100)/k